OpenStudy (mikey):

1/x^-3/6 help please i will give medal

alones (alones):

\[\frac{ 1 }{ x }^{\frac{- 3 }{ 6 }}\] this?

OpenStudy (3mar):

It will be like that

OpenStudy (3mar):

\[\frac{ 1 }{ x^{\frac{ -3 }{ 6 }} }\]

OpenStudy (mikey):

no its 1/x and the x has a square root of -3/6

OpenStudy (mikey):

yes 3mar

OpenStudy (3mar):

But the expression of AloneS works also, because one raised to any index equals one itself.

OpenStudy (3mar):

You need help, and I can help, what is the question?

OpenStudy (mikey):

what u showed in your second response

OpenStudy (3mar):

SO you got what you want? Hope that.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

First, reduce the fraction that is the exponent.

OpenStudy (mikey):

i have to rewrite in simplest radical form

OpenStudy (mathstudent55):

3/6 = ?

OpenStudy (mikey):

1/2

OpenStudy (mathstudent55):

Good. Now we have this: